# Intent3D: 3D Object Detection in RGB-D Scans Based on Human Intention

### v2

@@ Abstract
In real-life scenarios, humans seek out objects in the 3D world to fulfill their daily needs or intentions. This inspires us to introduce 3D intention grounding, a new task in 3D object detection employing RGB-D, based on human intention, such as "I want something to support my back". Closely related, 3D visual grounding focuses on understanding human reference. To achieve detection based on human intention, it relies on humans to observe the scene, reason out the target that aligns with their intention ("pillow" in this case), and finally provide a reference to the AI system, such as "A pillow on the couch". Instead, 3D intention grounding challenges AI agents to automatically observe, reason and detect the desired target solely based on human intention. To tackle this challenge, we introduce the new Intent3D dataset, consisting of 44,990 intention texts associated with 209 fine-grained classes from 1,042 scenes of the ScanNet [10] dataset. We also establish several baselines based on different language-based 3D object detection models on our benchmark. Finally, we propose IntentNet, our unique approach, designed to tackle this intention-based detection problem. It focuses on three key aspects: intention understanding, reasoning to identify object candidates, and cascaded adaptive learning that leverages the intrinsic priority logic of different losses for multiple objective optimization.
Figure 1: We introduce 3D intention grounding (right), a new task for detecting the object of interest using a 3D bounding box in a 3D scene, guided by human intention expressed in text (e.g., "I want something to support my back to relieve the pressure"). In contrast, the existing 3D visual grounding approach (left) relies on human reasoning and references for detection. The illustration clearly distinguishes that observation and reasoning are manually executed by human (left) and automated by AI (right).
@@ 1 Introduction
Performing object detection by following natural language instruction is crucial, as it can enhance the cooperation between humans and AI agents in real-world. Previous studies in Visual Grounding [23, 24, 25, 26, 42, 45, 56] have focused on localizing objects described by referential language in images. In recent years, research interests have expanded from Visual Grounding in 2D images to 3D indoor scenes, now recognized as 3D Visual Grounding (3D-VG) [6, 1, 52, 20, 4, 41, 60].
@@
In alignment with the sentiment expressed by Wayne Dyer, "Our intention creates our reality", humans identify and locate objects of interest for specific purposes or intentions. As depicted in Fig.1 (left), in real-world applications, to detect targets based on human intention, 3D-VG requires users to first observe the 3D world, reason about their desired target and provide a reference for the target’s category, attribution, or spatial relationship. However, various factors in daily use cases, such as engagement in intensive activities or experiencing visual impairments, may hinder humans from providing referential instructions. Therefore, the ability to automatically infer and detect the desired target based on user’s intentions, as shown in Fig.1 (right), becomes crucial and more intelligent.
@@
Recently, problem akin to reasoning on human instruction have been explored in 2D images, such as intention-oriented object detection [47], implicit instructions understanding [27], and affordance understanding [8, 40, 59, 39, 48]. However, their focus restricts to a partial view of our reality. Instead, 3D data mirrors the actual world more closely, featuring depth information, complete geometric and appearance characteristics of objects, and a comprehensive spatial context, elements that are lacking in 2D images, as highlighted in 3D object detection (3D-OD) [44, 54, 43, 38, 21, 49]. Furthermore, understanding 3D spaces plays a crucial role in the advancement of embodied AI [30, 13, 28], autonomous driving [19, 3, 9], and AR/VR applications [57, 31, 36]. Therefore, intention-based detection in 3D settings not only better simulates and addresses real-world complexities but also holds greater promise for future advancements.
@@
In this paper, we introduce the 3D Intention Grounding (3D-IG) task, which focuses on using human intention cues to automate reasoning and detection of desired objects in real-world 3D scenes. Firstly, we devise a data generation strategy to collect our 3D-IG dataset named Intent3D. Human intentions, being diverse, may face limitations in generalization with predefined formats. Hence, we used ChatGPT, enriched with human daily life knowledge, to generate intentions for each object. Our Intent3D dataset comprises 44,990 intention texts representing 209 fine-grained classes of objects from 1,042 ScanNet [10] point cloud scenes. Note that 3D-IG is a multi-instance detection problem, for example, given "I want to present slides to my coworkers", we can infer multiple instances of "Monitor" in Fig. 1, making 3D-IG even more challenging. Secondly, to fully gauge our current research’s capability in solving this problem, we construct several baselines using three main language-based 3D-OD techniques for our benchmark. This involves evaluating our dataset using the following models: state-of-the-art expert model specifically designed for 3D visual grounding, existing foundation model formulated for generic 3D understanding tasks, and Large Language Model (LLM)-based model. We evaluate these baselines with multiple settings, namely training from scratch, fine-tuning, and zero-shot. Thirdly, we propose IntentNet, our novel method, to tackle this intention-based 3D object detection problem. To achieve a comprehensive understanding of intention texts, we introduce the Verb-Object Alignment technique, focusing on initially recognizing and aligning with the verb and subsequently aligning with the corresponding object features. For explicitly identifying relevant object candidates in the sparse 3D point cloud, we incorporate the Candidate Box Matching component. Furthermore, to concurrently optimize the intention understanding, candidate box matching, and detection objectives, we propose a novel Cascaded Adaptive Learning mechanism to adaptively scale up different losses based on their intrinsic priority logic. We summarize our contributions in this paper as follows:
@@
  • We introduce a new task, 3D Intention Grounding (3D-IG), which takes a 3D scene and free-form text describing the human intention as input to directly detect the target of interest. To construct a benchmark for the 3D-IG task, we compile a new dataset called Intent3D, comprising 44,990 intention texts covering 209 fine-grained classes of objects from 1,042 3D scenes.
  • We construct comprehensive baselines on our Intent3D, encompassing three main classes of language-based 3D-OD methods: expert model, foundation model, and LLM-based model.
  • We propose a new method, IntentNet, for intention-based detection.
  • We plan to release the dataset, source code and model checkpoints upon acceptance of the paper for future research development.
@@ 2 Related work
Referential Language Grounding Visual Grounding (VG), also known as Referring Expression Comprehension [23, 24, 25, 26, 42, 45, 56], aims to detect the object by a bounding box, given a 2D image and a human referential instruction pertaining to the object of interest. Utilizing point cloud scans from ScanNet [10], recently, ScanRefer [6] extended this task to 3D Visual Grounding (3D-VG), aiming to predict a 3D bounding box for the target object given a 3D point cloud of the indoor environment and a human referential language. Similarly, ReferIt3D [1] proposes the 3D-VG task also based on ScanNet [10] but with a simpler configuration, where it supplies ground truth boxes for all candidate objects in the scene and only needs approaches to choose the correct bounding box. Nevertheless, existing 3D-VG benchmarks predominantly evaluate the capability of methods in interpreting human referential instructions, typically encompassing the category, attribution, or spatial relationship of the object. The exploration of directly inferring and detecting the desired target based on human intention instructions in the 3D domain remains an uncharted area.
@@
Affordance Understanding Derived from the concept in [14], affordance detection aims to localize an object or a part of object in 2D [8, 40, 39, 48] or 3D [59, 11] focusing on feasible/affordable actions that can be executed while engaging with the object of interest. However, the affordable action still cannot convey human intention; for example, the action "grasp a bottle" does not necessarily equate to the underlying intention "I want something to drink". Furthermore, the dependency on phrase expressions [39], closed-set classifications [8, 40, 59, 48], and object-centered approaches [11] restricts its ability to interpret free-form intention and make inferences about 3D scenes.
@@
Complex Instruction Understanding To explore more complex instruction understanding, RIO [47] introduces a dataset and solutions for intention-oriented object detection based on MSCOCO [33] 2D images. Moreover, previous studies also utilize the advancement of large multimodal models to address instruction reasoning in 2D open environments. LISA [27] adopts large language model for understanding complex and implicit query texts to generate segmentation masks. Ferret [55] reasons about multiple targets involved in instructions and introduces a dataset enriched with spatial knowledge to enhance model robustness. However, reasoning in 3D environments based on complex instructions, especially human intentions, remains underexplored.
@@
Language-based 3D Object Detection Methods Currently, the 3D-VG community has developed various expert (specialist) models [6, 1, 52, 20, 4, 41, 60], trained from scratch to comprehend the free-form language and sparse point clouds for detection task. State-of-the-art expert models like BUTD-DETR [20] and EDA [52] inherit the merit from MDETR [22], focusing on understanding the target’s category or attribution expressed in the language. 3D-VisTA [62], on the other hand, introduces a transformer-based foundation model to jointly train across multiple 3D tasks, including 3D-VG, 3D Dense captioning, 3D Question Answering, and 3D Situated Reasoning. Furthermore, inspired by advancements made in large multimodal models [2, 29, 35, 34, 50, 58], researchers are actively exploring comprehension of point cloud information through the utilization of large language models. Models such as PointLLM [53] and Point-Bind [15] primarily focus on object point cloud scans, while 3D-LLM [16], Chat-3D [51], and Chat-3D v2[18] are designed to comprehend complex 3D environment point cloud scans and perform detection.
@@ 3.1 Task Definition
3D Intention Grounding aims to detect 3D objects within a 3D scene to match a given human intention. It takes as input a 3D scene (e.g. "living room"), represented by a point cloud, and a free-form text (e.g. "I need a place to sit down and relax") describing a human intention. The goal is to predict all the 3D bounding boxes for all instances of the target object (i.e. detect all the "chair").
Figure 2: (Upper row) Flowchart for dataset construction. Using ScanNet’s data, we first select objects by three principal criteria: Common Object, Non-trivial Object, Unambiguous object. Then, we employ ChatGPT to generate intention texts given the prompt we designed. Finally, we manually clean the data. (Lower row) Examples from our dataset for different number of targets and length of texts.
@@ 3.2 Intent3D Dataset
Our 3D scenes and bounding box annotations are sourced from ScanNet [10], which includes 1,513 indoor environment scenes of point clouds. We use ChatGPT (GPT-4) to generate human intention text for selected objects in the 3D scenes. The pipeline for our dataset construction is shown in Fig. 2 (upper row).
Step 1: Object Selection
@@
We select the objects in each scene based on following three principal criteria. 1) Common Objects: We focus on common indoor objects with diverse human intentions, excluding structural elements like "wall" and "ceiling", and opting for objects occurring in at least four distinct scenes. 2) Non-trivial Objects: In each scene, we select objects with fewer than six instances in their fine-grained categories, thus excluding overly apparent objects and enhancing the challenge of our grounding task. 3) Unambiguous objects: Addressing ambiguity is crucial in generating human intention text, as objects with different fine-grained classes can fulfill the same human intention due to shared higher-level categories (e.g., "floor lamp" and "desk lamp" both classified under "lamp"), synonyms (e.g., "trash bin" vs. "garbage bin"), and variations between singular and plural terms (e.g., "bookshelf" vs. "bookshelves"). Furthermore, ambiguity also arises when objects belong to different higher-level categories but can fulfill the same human intention (e.g., both "TV" and "Monitor" can satisfy the intention "I want to display my chart to my co-workers"). To mitigate this, we manually identify ambiguous situations for each scene and filter out objects that serve similar human intentions.
Step 2: Text Generation
@@
Unlike referential language in 3D-VG, which is associated with the deterministic characteristic of objects, intention language in 3D-IG stems from the diverse and spontaneous nature of human daily needs. Hence, instead of prescribing specific formats or contents [6, 1] for language generation, we leverage ChatGPT (GPT-4), endowed with extensive real-world knowledge, to craft anthropomorphic intention text using a question-and-answer method. We construct the question in two parts: one for constraints (Prompt-1) and the other for scene context (Prompt-2). Specifically, the first part is designed to emulate human daily expressions of their intentions and preventing the disclosure of any referential details, such as the object’s category or location, in the response. This is done to challenge AI agents to engage in reasoning independently:
@@
Prompt-1: "You are a helpful assistant in providing human intention towards each object. The intention is from a first-person perspective, in the format of ‘I want / need / intend / … to …’. The intention must avoid mentioning synonyms, categories, locations, or attributes of the object."
@@
In the second part, we supply ChatGPT with relevant information about the scene using the following template:
Prompt-2: “In a $<$scene$>$ with $<$objects$>$, what can you do with each object?"
@@
where <$scene$> refers to the type of scene and <$objects$> implies a list of all object classes present in the scene. On average, we generate six intention texts per object from ChatGPT. Fig. 2 (lower row) displays some samples we generated with varying lengths and different numbers of targets.
Step 3: Data Cleaning
@@
Due to occasional nonsensical responses from ChatGPT, we manually filter out gibberish from the generated text. Furthermore, we discard responses that fail to generate. In cases of ambiguity, such as with the term "keyboard" in our dataset, where it should refer to a computer keyboard rather than a musical instrument, we manually recreate the correct intention sentences. Additionally, in samples with repeated intention sentences, we manually regenerate them to ensure diversity.

### v3

@@ Abstract
In real-life scenarios, humans seek out objects in the 3D world to fulfill their daily needs or intentions. This inspires us to introduce 3D intention grounding, a new task in 3D object detection employing RGB-D, based on human intention, such as “I want something to support my back.” Closely related, 3D visual grounding focuses on understanding human reference. To achieve detection based on human intention, it relies on humans to observe the scene, reason out the target that aligns with their intention (“pillow” in this case), and finally provide a reference to the AI system, such as “A pillow on the couch”. Instead, 3D intention grounding challenges AI agents to automatically observe, reason and detect the desired target solely based on human intention. To tackle this challenge, we introduce the new Intent3D dataset, consisting of 44,990 intention texts associated with 209 fine-grained classes from 1,042 scenes of the ScanNet [Dai et al., 2017] dataset. We also establish several baselines based on different language-based 3D object detection models on our benchmark. Finally, we propose IntentNet, our unique approach, designed to tackle this intention-based detection problem. It focuses on three key aspects: intention understanding, reasoning to identify object candidates, and cascaded adaptive learning that leverages the intrinsic priority logic of different losses for multiple objective optimization.
@@ 1 Introduction
Performing object detection by following natural language instruction is crucial, as it can enhance the cooperation between humans and AI agents in real-world. Previous studies in Visual Grounding [Kang et al., ; Kang et al., 2024a; b; Kazemzadeh et al., 2014; Mao et al., 2016; Plummer et al., 2015; Yu et al., 2016] have focused on localizing objects described by referential language in images. In recent years, research interests have expanded from Visual Grounding in 2D images to 3D indoor scenes, now recognized as 3D Visual Grounding (3D-VG) [Chen et al., 2020; Achlioptas et al., 2020; Wu et al., 2023; Jain et al., 2022; Cai et al., 2022; Luo et al., 2022; Zhao et al., 2021].
@@
In alignment with the sentiment expressed by Wayne Dyer, “Our intention creates our reality”, humans often identify and locate objects of interest for specific purposes or intentions. A recent demonstration by OpenAI featuring a humanoid robot [OpenAI, 2024] showcases the promising potential of intention-based understanding. As depicted in Fig.1 (left), in the real-world applications, to detect targets based on human intention, 3D-VG requires users to first observe the 3D world, reason about their desired target and provide a reference for the target’s category, attributes, or spatial relationship. However, various factors in daily use cases, such as engagement in intensive activities or experiencing visual impairments, may hinder humans from providing referential instructions. Therefore, the ability to automatically infer and detect the desired target based on the user’s intentions, as shown in Fig.1 (right), becomes crucial and more intelligent.
@@
Recently, problems akin to reasoning on human instruction have been explored in 2D images, such as intention-oriented object detection [Qu et al., 2024], implicit instructions understanding [Lai et al., 2023], and affordance understanding [Chuang et al., 2018; Luo et al., 2021; Zhai et al., 2022; Lu et al., 2022; Sawatzky et al., 2019]. However, their focus is restricted to a partial view of our reality. Instead, 3D data mirrors the actual world more closely, featuring depth information, complete geometric and appearance characteristics of objects, and a comprehensive spatial context, elements that are lacking in 2D images, as highlighted in 3D object detection (3D-OD) [Pan et al., 2021; Yang et al., 2018; Misra et al., 2021; Liu et al., 2021b; Jiang et al., 2020; Schult et al., 2022]. Furthermore, understanding 3D spaces plays a crucial role in the advancement of embodied AI [Li et al., 2023b; Gao et al., 2023; Li et al., 2023a], autonomous driving [Huang et al., 2018; Caesar et al., 2020; Cui et al., 2021], and AR/VR applications [Yu et al., 2022; Lim et al., 2022; Liu et al., 2021a]. Therefore, intention-based detection in 3D settings not only better simulates and addresses real-world complexities but also holds greater promise for future advancements.
Figure 1: We introduce 3D intention grounding (right), a new task for detecting the object of interest using a 3D bounding box in a 3D scene, guided by human intention expressed in text (e.g., “I want something to support my back to relieve the pressure”). In contrast, the existing 3D visual grounding (left) relies on human reasoning and references for detection. The illustration clearly distinguishes that observation and reasoning are manually executed by human (left) and automated by AI (right).
@@
In this paper, we introduce the 3D Intention Grounding (3D-IG) task, which focuses on using human intention cues to automate reasoning and detection of desired objects in real-world 3D scenes. Firstly, we devise a data generation strategy to collect our 3D-IG dataset, named Intent3D. Human intentions are diverse. Therefore, it limits generalization if we use predefined formats for human annotators to label the intention sentences. Moreover, as highlighted in [Yang et al., 2024; Abdelreheem et al., 2024; Ding et al., 2022; Tan et al., 2024], crowd-sourced annotators often lack reliability, while professional annotators are expensive and not scalable for large-scale development. Instead, human quality checks [Yang et al., 2024] show that LLMs (GPT-4 [OpenAI, ]) can serve as human-level annotators, achieving a low error rate. Given these two reasons, we adopt ChatGPT, enriched with diverse daily life knowledge, to generate intentions for each object. Our Intent3D dataset comprises 44,990 intention texts representing 209 fine-grained classes of objects from 1,042 ScanNet [Dai et al., 2017] point cloud scenes. Note that 3D-IG is a multi-instance detection problem. For example, given “I want to present slides to my coworkers,” we can infer multiple instances of “Monitor” in Fig. 1 (right part), making 3D-IG even more challenging. Secondly, to fully gauge our current research capability in solving this problem, we construct several baselines using three main language-based 3D-OD techniques for our benchmark. This involves evaluating our dataset using the following models: expert models specifically designed for 3D visual grounding, an existing foundation model formulated for generic 3D understanding tasks, and a Large Language Model (LLM)-based model. We evaluate these baselines with multiple settings, namely training from scratch, fine-tuning, and zero-shot. Lastly, we propose IntentNet, a novel method to tackle the 3D-IG problem. To achieve a comprehensive understanding of intention texts, we introduce the Verb-Object Alignment technique, which focuses on initially recognizing and aligning with the verb and subsequently aligning with the corresponding object features. We further propose the Candidate Box Matching component to explicitly identify relevant object candidates in the sparse 3D point cloud. To concurrently optimize the intention understanding, candidate box matching, and detection objectives, we incorporate a novel Cascaded Adaptive Learning mechanism to adaptively scale up different losses based on their intrinsic priority logic.
@@
In summary, our contributions are threefold: (i) We introduce a new task, 3D Intention Grounding (3D-IG), which takes a 3D scene and a free-form text describing the human intention as input to directly detect the target of interest. To construct a benchmark for the 3D-IG task, we compile a new dataset called Intent3D, comprising 44,990 intention texts covering 209 fine-grained classes of objects from 1,042 3D scenes. (ii) We construct comprehensive baselines for our Intent3D, encompassing three main classes of language-based 3D-OD methods: expert model, foundation model, and LLM-based model. (iii) We propose a novel method, IntentNet, for intention-based detection, achieving state-of-the-art performance on the Intent3D benchmark.
@@ 2 Related work
Referential Language Grounding Visual Grounding (VG)  [Kang et al., ; Kang et al., 2024a; b; Kazemzadeh et al., 2014; Mao et al., 2016; Plummer et al., 2015; Yu et al., 2016] aims to detect the object by a bounding box, given a 2D image and a human referential instruction pertaining to the object of interest. Utilizing point cloud scans from ScanNet [Dai et al., 2017], ScanRefer [Chen et al., 2020] recently extended this task to 3D Visual Grounding (3D-VG), aiming to predict a 3D bounding box for the target object given a 3D point cloud of the indoor environment and a human referential language instruction. Similarly, ReferIt3D [Achlioptas et al., 2020] proposes the 3D-VG task also based on ScanNet [Dai et al., 2017] but with a simpler configuration, where it supplies ground truth boxes for all candidate objects in the scene and only needs method to choose the correct bounding box. Nevertheless, existing 3D-VG benchmarks predominantly evaluate the capability of methods in interpreting human referential instructions, typically encompassing the category, attributes, or spatial relationship of the object. The exploration of directly inferring and detecting the desired target based on human intention instructions in the 3D domain remains an uncharted area.
@@
Affordance Understanding Derived from the concept in [Gibson, 2014], affordance detection aims to localize an object or a part of object in 2D [Chuang et al., 2018; Luo et al., 2021; Lu et al., 2022; Sawatzky et al., 2019] or 3D [Zhai et al., 2022; Deng et al., 2021] focusing on feasible/affordable actions that can be executed while engaging with the object of interest. However, the affordable action still cannot convey human intention; for example, the action “grasp a bottle” does not necessarily equate to the underlying intention “I want something to drink”. Furthermore, the dependency on phrase expressions [Lu et al., 2022], closed-set classifications [Chuang et al., 2018; Luo et al., 2021; Zhai et al., 2022; Sawatzky et al., 2019], and object-centered approaches [Deng et al., 2021] restricts the ability of affordance detection to interpret free-form intention and infer 3D scenes.
@@
Complex Instruction Understanding To explore more complex instruction understanding, RIO [Qu et al., 2024] introduces a dataset for intention-oriented object detection based on MSCOCO [Lin et al., 2014] 2D images. Moreover, previous studies have also utilized large language models (LLMs) to address instruction reasoning in 2D images. LISA [Lai et al., 2023] adopts LLM to understand complex and implicit query texts. Ferret [You et al., 2023] reasons about multiple targets involved in instructions and introduces a dataset enriched with spatial knowledge. However, reasoning in 3D scene based on human intentions, remains underexplored.
@@
Language-based 3D Object Detection Methods Currently, the 3D-VG community has developed various expert (specialist) models [Chen et al., 2020; Achlioptas et al., 2020; Wu et al., 2023; Jain et al., 2022; Cai et al., 2022; Luo et al., 2022; Zhao et al., 2021], trained from scratch to comprehend the free-form language and sparse point clouds for detection task. State-of-the-art expert models like BUTD-DETR [Jain et al., 2022] and EDA [Wu et al., 2023] inherit the merits from MDETR [Kamath et al., 2021], focusing on understanding the target’s category or attributes expressed in the language. 3D-VisTA [Zhu et al., 2023], on the other hand, introduces a transformer-based foundation model to jointly train across multiple 3D tasks. Furthermore, inspired by advancements made in LLMs [Liu et al., 2023; Shang et al., 2024; Yuan et al., 2024], researchers have also explored the utilization of LLMs in 3D domain. Models such as PointLLM [Xu et al., 2023] and Point-Bind [Guo et al., 2023] primarily focus on object point cloud scans, while 3D-LLM [Hong et al., 2024], Chat-3D [Wang et al., 2023], and Chat-3D v2[Huang et al., 2023] are designed for 3D environment scans. We comprehensively evaluate these methods on our Intent3D benchmark. Unlike previous approaches, our proposed IntentNet explicitly models intention-based multimodal reasoning.
@@ 3.1 Task Definition
3D Intention Grounding aims to detect 3D objects within a 3D scene to match a given human intention. It takes as input a 3D scene (e.g. “living room”), represented by a point cloud, and free-form text (e.g. “I need a place to sit down and relax”) describing a human intention. The goal is to predict all the 3D bounding boxes for all instances of the target object (i.e. detect all the “chairs”).
Figure 2: (Upper row) Flowchart for dataset construction. After constructing the scene graph, we select objects by three criteria: Common Object, Non-trivial Object, Unambiguous Object. We use ChatGPT to generate intention texts given the prompt we designed. Finally, we manually clean the data. (Lower row) Examples in our dataset for different number of targets and length of texts.
@@ 3.2 Intent3D Dataset
The pipeline for our dataset construction is shown in Fig. 2 (upper row).
Step 1: Scene Graph Construction
First of all, given a 3D point cloud of a scene, we construct a scene graph as a json file to organize all the annotations or information, including the type of the scene, the category at different levels of granularity for each object within the scene, the number of instances sharing the same category, and the 3D bounding box (x, y, z, w, l, h) for each object.
Step 2: Object Selection
@@
We select the objects in each scene based on the following three criteria. 1) Common Objects: We focus on common indoor objects with diverse human intentions, excluding structural elements like “wall” and “ceiling”, and opting for objects occurring in at least four distinct scenes. 2) Non-trivial Objects: In each scene, we select objects with fewer than six instances in their fine-grained categories, thus excluding overly apparent objects and enhancing the challenge of our grounding task. For example, in a conference room, we select the “table” with only one instance rather than the “chair” with ten instances. 3) Unambiguous objects: Addressing ambiguity is crucial in generating human intention text, as objects with different fine-grained classes can fulfill the same human intention due to shared higher-level categories (e.g., “floor lamp” and “desk lamp” both classified under “lamp”), synonyms (e.g., “trash bin” vs. “garbage bin”), and variations between singular and plural terms (e.g., “bookshelf” vs. “bookshelves”). Furthermore, ambiguity also arises when objects belong to different higher-level categories but can fulfill the same human intention (e.g., both “TV” and “Monitor” can satisfy the intention “I want to display my chart to my co-workers”). To mitigate this, we manually identify ambiguous situations and filter out objects that serve similar human intentions.
Step 3: Text Generation
@@
We leverage ChatGPT (GPT-4) to generate intention texts in a question-answering manner. The question has two parts: one for constraints (Prompt-1) and the other for scene context (Prompt-2). Specifically, the first part is designed to emulate human expressions of their intentions and prevent the disclosure of referential details, such as the object’s category or location, in the response. This is done to challenge AI agents to engage in reasoning independently:
@@
Prompt-1: “You are a helpful assistant in providing human intention towards each object. The intention is from a first-person perspective, in the format of ‘I want / need / intend / … to …’. The intention must avoid mentioning synonyms, categories, locations, or attributes of the object.
@@
In the second part, we supply ChatGPT with relevant information about the scene:
Prompt-2: “In a $<$scene$>$ with $<$objects$>$, what can you do with each object?
@@
where $<$scene$>$ refers to the type of scene and $<$objects$>$ implies a list of all object classes present in the scene. On average, we generate six intention texts per object. Fig. 2 (lower row) displays some samples we generated with varying lengths and different numbers of targets.
Step 4: Data Cleaning
@@
Due to occasional nonsensical responses from ChatGPT, we manually filter out gibberish from the generated text. Furthermore, we discard responses that fail to generate. In cases of ambiguity, such as “keyboard” in our dataset, where it should refer to a computer keyboard rather than a musical instrument, we manually recreate the sentences. Additionally, in samples with repeated intention sentences, we manually regenerate them to ensure diversity. In our Appendix A, we provide more details (A.1) of failed responses from ChatGPT and the usage of API, and we further discuss ambiguous problem (A.4) and human-likeness test (A.5) of our intention sentences.